# Scaling Linear Attention with Sparse State Expansion

### v1

@@ 6 Limitations and Future Work
While our current experiments demonstrate SSE’s promising scaling performance with GLA-style transition mechanisms, we did not find significantly improved retrieval over the baseline in experiments using delta-rule mechanisms (e.g., GDN). We hypothesize this is because non-diagonal transition matrices, which involve rapid inter-row interactions and fast information dynamics at every step, challenge simple classifiers to fit effectively. We believe more sophisticated classification functions, such as state-aware classifiers, could effectively address these issues, representing a promising avenue for future research.
@@
We acknowledge that SSE’s current context-only hard-selection mechanism may not be optimal and warrants broader exploration. A compelling future direction involves incorporating dependencies on positional and historical information, leading to a more general form such as $\mathbf{e}_{t}=g(t,\mathbf{x}_{t},\mathbf{S}_{t-1})$.
@@ 7 Related Work
State Expansion. Recent literature [2, 3, 23, 1, 49] consistently highlights the challenges posed by limited state size in linear attention models, particularly evident in retrieval tasks. Consequently, state expansion has emerged as a seemingly indispensable direction. Beyond directly expanding the state, Log-Linear Attention [21] adopts a Fenwick tree structure to enable logarithmically growing state sizes. Our work bears the closest resemblance to Mixture-of-Memories (MoM) [13], which models the state using a Mixture-of-Experts (MoE) approach and shares a similar sparse spirit with SSE. However, SSE designs a sparse state expansion scheme specifically from an information classification perspective, while also decoupling the influence of parameter count. Furthermore, we provide a comprehensive theoretical framework demonstrating that SSE’s state achieves more effective information storage. Our extensive experimental pipeline, spanning pretraining to RL, robustly confirms the proposed model’s stability.
@@ 8 Conclusion
In this work, we address the limitations of existing linear attention models in handling long contexts by proposing two key innovations: row-sparse state updates and Sparse State Expansion (SSE). Our framework conceptualizes state updates as an information classification process and enables efficient expansion of contextual states into multiple partitions with controlled parameter count. This design facilitates both scalable memory capacity and more discriminative state representations. We demonstrate that SSE and its hybrid variant (SSE-H) achieve superior performance across language modeling, in-context retrieval, and mathematical reasoning tasks. Notably, our 2B SSE-H model sets state-of-the-art reasoning performance among small-scale models on math-focused benchmarks, significantly surpassing comparably sized Transformers. Together with its favorable scalability, these results position SSE as a promising and efficient architecture for high-fidelity long-context modeling.
@@ 9 Contributions
Yuqi Pan1,$2$Work done during internship at ByteDance Seed., Yongqi An1,$2$Work done during internship at ByteDance Seed., Zheng Li1

### v2

@@ 6 Limitations and Future Work
While our experiments demonstrate the promising scalability of SSE with GLA-style transition mechanisms, we have not yet conducted comprehensive evaluations with delta-rule baselines (e.g., GDN). Moreover, more sophisticated classification functions, such as state-aware classifiers, could be effectively combined with SSE-GDN, offering a promising direction for future work.
@@
We acknowledge that SSE’s current context-only hard-selection mechanism may not be optimal and warrants broader exploration. The computation of $\mathbf{e}_{t}$ provides a flexible interface for incorporating inductive biases into the model. A compelling future direction involves incorporating dependencies on positional and historical information, leading to a more general form such as $\mathbf{e}_{t}=g(t,\mathbf{x}_{t},\mathbf{S}_{t-1})$.
@@ 7 Related Work
State Expansion. Recent literature [2, 3, 23, 1, 49] consistently highlights the challenges posed by limited state size in linear attention models, particularly evident in retrieval tasks. Consequently, state expansion has emerged as a seemingly indispensable direction. Beyond directly expanding the state, Log-Linear Attention [21] adopts a Fenwick tree structure to enable logarithmically growing state sizes. Our work bears the closest resemblance to Mixture-of-Memories (MoM) [13], which models the state using a Mixture-of-Experts (MoE) approach and shares a similar sparse spirit with SSE. However, SSE designs a sparse state expansion scheme specifically from an information classification perspective, explicitly decouples state capacity from parameter count, and introduces improved feature mapping and gating mechanisms. Furthermore, we provide a comprehensive theoretical framework demonstrating that SSE’s state achieves more effective information storage. Our extensive experimental pipeline, spanning pretraining to RL, robustly confirms the proposed model’s stability.
@@ 8 Conclusion
In this work, we enhance the capacity of existing linear attention models in handling long contexts by proposing two key innovations: row-sparse state updates and Sparse State Expansion (SSE). Our framework conceptualizes state updates as an information classification process and enables efficient expansion of contextual states into multiple partitions with controlled parameter count. This design facilitates both scalable memory capacity and more discriminative state representations. We demonstrate that SSE and its hybrid variant (SSE-H) achieve superior performance across language modeling, in-context retrieval, and mathematical reasoning tasks. Notably, our 2B SSE-H model sets state-of-the-art reasoning performance among small-scale models on math-focused benchmarks, significantly surpassing comparably sized Transformers. Together with its favorable scalability, these results position SSE as a promising and efficient architecture for high-fidelity long-context modeling.
@@ 9 Contributions
Yuqi Pan1,2222Work done during internship at ByteDance Seed., Yongqi An1,2222Work done during internship at ByteDance Seed., Zheng Li1